sxdsouls:

A car jack with a mechanical advantage of 6 needs to produce an output force of 600 N to raise a car. What input force is required? 6 N 100 N 600 N 3,600 N

Astrid1:

Input force = 100N Explanation: Formula Mechanical Advantage = output force / input force. Givens MA = 6 Output Force = 600 N Input Force = ?? Solution 6 = 600 / x Multiply both sides by x 6x = (600/x)*x 6x = 600 The xs cancel. Now Divide by 6 6x/6 = 600/6 x = 100 N (Found on allanswers.com)

sxdsouls:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Astrid1 Input force = 100N ty Explanation: Formula Mechanical Advantage = output force / input force. Givens MA = 6 Output Force = 600 N Input Force = ?? Solution 6 = 600 / x Multiply both sides by x 6x = (600/x)*x 6x = 600 The xs cancel. Now Divide by 6 6x/6 = 600/6 x = 100 N (Found on allanswers.com) \(\color{#0cbb34}{\text{End of Quote}}\)

Astrid1:

I hope it helps you even if it was found off the internet.